OpenStudy (anonymous):

A small college had 2546 students in 2002 and 2702 students in 2004. If the enrollment follows a linear growth pattern, How many students will the collage have in 2008?

OpenStudy (chaise):

Let y = amount of children. Let x = years past Let c = amount of children in 2002. m = (y-y)/(x-x) m = (2702-2546)/(2004-2002) m = 78 y = mx + c y = 78x + c your c is always your initial value. SO: y = 78x + 2546 So basicly this graph tells us that each year there are 78 people coming to the school. 6 years will have past since 2002. So we just change x to 6. y = 78*6 + 2546. y = 3014 Does this make sense?

OpenStudy (chaise):

x should really be years since 2002. Sorry.

OpenStudy (anonymous):

thank you very much I have more question. I hope you help me for those too.

OpenStudy (chaise):

I can do my best. shoot :)